# A complex scalar field theory for charged fluids, superfluids, and fracton fluids

### v1

@@ 1 Introduction
Fluid dynamics is one of the oldest and most established fields of physics, with applicability ranging from microscopic quantum fluids, such as the quark-gluon plasma, all the way to the universe itself. Despite its long and illustrious history, a complete understanding of fluids is still lacking. Some of the most pressing open problems include nonlinear phenomena such as turbulence and the million-dollar problem on the Navier–Stokes equations, but also the behavior of strongly correlated quantum fluids, and questions concerning their existence (or lack thereof) at zero temperature Endlich:2010hf ; Dersy:2022kjd ; Cuomo:2024ekf .
@@
From a theoretical standpoint, a more compelling picture is offered by a field-theoretic formulation of hydrodynamics in terms of a local action Son:2002zn ; Dubovsky:2005xd ; Dubovsky:2011sj ; Nicolis:2011cs ; Kovtun:2014hpa ; Crossley:2015evo . In this framework, the hydrodynamic equations follow from the Euler–Lagrange equations for the relevant low-energy fields, while the path-integral formalism grants access to the correlation functions and scattering amplitudes (see, e.g., PhysRevLett.94.175301 ; Endlich:2010hf ; Nicolis:2011cs ; Dersy:2022kjd ; Jain:2023obu ; Delacretaz:2023ypv ; Cuomo:2024ekf ). In short, the action formulation allows the use of the full quantum field theory (QFT) toolkit. The models in question are effective field theories (EFTs) where infrared degrees of freedom are the Nambu-Goldstone bosons of spontaneously broken symmetries. Unfortunately, without microscopic input from the UV, these EFTs can suffer from various pathologies.
@@
where $D_{\mu}\psi=\partial_{\mu}\psi-A_{\mu}$ is a covariant derivative and $\mathcal{L}(X)$ is an arbitrary function. For example, choosing $\mathcal{L}(X)=X$ gives a simple quadratic theory with linear equations of motion. In general, however, the resulting equations are nonlinear and lead to the formation of caustic singularities in finite time, wherein the characteristics intersect and second-order derivatives diverge, signaling the breakdown of the EFT Babichev:2016hys . These issues can be circumvented by introducing a UV completion in terms of a complex scalar field, which reproduces the same low-energy dynamics as the EFT model (up to the formation of caustic singularities in the EFT) while regularizing the caustic instability Babichev:2017lrx .
For superfluids, a suitable UV completion is given by a canonical complex scalar field action Babichev:2018twg
@@
where $D_{\mu}\Phi=\left(\partial_{\mu}-iA_{\mu}\right)\Phi$. By placing the system at finite chemical potential via a constant background gauge field $A_{\mu}=(\mu_{0},0)$ with $\mu_{0}>m$, and implementing the polar decomposition $\Phi=\sqrt{\rho}e^{\psi}$, the theory Eq. (2) reproduces Eq. (1) in the long-wavelength limit after integrating out the Higgs mode Joyce:2022ydd . Clearly, the theory Eq. (2) contains more information than the EFT Eq. (1). For example, it can be used to compute the healing length of the superfluid. It is precisely the absence of information about the short-range physics in the EFT Eq. (1) that manifests itself in the appearance of caustic singularities.
Can it happen that some of the long-standing problems in hydrodynamics can similarly be addressed by invoking short-range physics?
Driven by these questions, in this manuscript, we propose a UV completion to the EFTs of charged fluids Dubovsky:2011sj and superfluids at finite-temperature Nicolis:2011cs . For this purpose, we utilize a geometric picture of fluid flow in terms of a comoving spatial hypersurface and define on it a complex scalar field transforming in the linear representation of the $U(1)_{Q}$ charge symmetry. The models proposed herein reproduce the established constitutive relations for the ideal part of hydrodynamic currents, and the hydrodynamic excitations are identified with the Nambu-Goldstone modes corresponding to small fluctuations around the state of thermodynamic equilibrium.
We show that ordinary charged fluids are necessarily invariant under a restrictive chemical shift symmetry that freezes the dynamics of the elementary charges in the comoving space. In this interpretation, charges exhibit fractonic immobility on the comoving plane but are dragged through spacetime by the fluid flow. This symmetry is relaxed in the superfluid phase in order to account for the dynamics of the superfluid component. Our formulation naturally accommodates a classification of exotic fluid phases interpolating between normal fluids and superfluids. These fracton fluids exhibit partially restricted mobility, a consequence of the comoving multipole symmetries [^1] Analogous ”crystal-multipole” symmetries were introduced in Jain:2024ngx ., which in our approach emerge as a finite truncation of the chemical shift symmetries.
@@ 2 Comoving formulation of fluid dynamics
In this section we review the comoving formulation of perfect fluids and encode their relativistic hydrodynamics in an action principle. The variational approach to fluid mechanics in terms of a set of comoving coordinates has a long history 10.1063/1.1706053 ; annurev:/content/journals/10.1146/annurev.fl.20.010188.001301 ; Bahcall:1991an , see also jackiw2002lectures for a pedagogical introduction. More recently, the problem has been revisited from a modern effective field theory viewpoint Dubovsky:2005xd , which we will mostly follow in this section. We also provide a complementary geometrical interpretation of the comoving formulation in terms of spacelike hypersurface embedded in the physical spacetime. Our setup aligns closely with the geometric framework of elasticity theory introduced in Armas:2019sbe . Although the comoving hypersurface perspective may currently appear as a mere reinterpretation, it sets the stage for the constructions developed in the upcoming sections.
@@ 2.3 Effective action for uncharged fluids
Therefore, the leading order theory for the hydrodynamics of uncharged fluids can be expressed as Dubovsky:2005xd
@@
There are also conservation laws associated with $\text{SDiff}(\mathbb{R}^{2})$ symmetry Eq. (6), corresponding to vorticity conservation (see Cuomo:2024ekf ).
@@ 3 Charged fluids
Utilizing the comoving hypersurface interpretation, we propose a field-theoretic model for charged fluid dynamics that reproduces the nonlinear hydrodynamic equations of a perfect fluid. Contrary to previous formulations Dubovsky:2011sj , the charge symmetry is realized linearly on the complex scalar field $\Phi$ residing on the comoving hypersurface. We provide a first-principles proof for the emergence of the chemical shift symmetry in charged fluids, corresponding to the comoving conservation of all multipole moments, and elucidate the resulting symmetry structure. Then, we formulate a symmetry-invariant action, derive the constitutive relations, and establish an explicit dictionary mapping the fields of the theory to hydrodynamic variables. We also discuss the equilibrium state, its symmetry-breaking pattern, and the spectrum of Nambu–Goldstone modes.
@@ 3.1 Multipole symmetries on the comoving plane
These charges reflect the conservation of the multipole moments of the charge density, familiar in the context of fractons Gromov:2018nbv , with respect to the comoving hypersurface. Such conservation laws impose strong constraints on the dynamics of the elementary charges giving rise to fractonic behavior, wherein charges cannot move along the comoving hypersurface without violating one of the higher-moment conservation laws [^5] Nevertheless, they can still be advected through spacetime by a fluid’s motion.. A theory of charged fluids must respect the conservation of the multipole moments Eq. (38) and therefore be invariant under the symmetries generated by these charges,
@@
where $\Lambda_{I_{1}\dots I_{N}}$ is a constant symmetric tensor of parameters. Since this must hold for any $N$, we conclude that in ordinary fluids the charge symmetry Eq. (34) must be promoted to the chemical shift symmetry [^6] The chemical shift symmetry was first introduced in Dubovsky:2011sj to construct an effective field theory for hydrodynamics. In our formulation, however, it is realized linearly on the complex scalar field and emerges naturally as a consequence of Noether’s theorem.
@@ 3.2 The model
Having established the appropriate degrees of freedom for charged fluids and their symmetry structure, we are now equipped to write down an action governing their hydrodynamics. At leading order in derivatives, the most general such action takes the form [^7] The theory Eq. (48) can be understood as a UV completion of the effective hydrodynamic theory constructed in Dubovsky:2011sj .
@@
where $D_{0}\Phi=u^{\mu}(\partial_{\mu}-iA_{\mu})\Phi$ and $E(|\Phi|,b)$ is an arbitrary function. Throughout this manuscript, we assume that $E(|\Phi|,b)$ is factorizable [^8] This choice corresponds to a subclass of fluids that admit a factorizable equation of state., $E(|\Phi|,b)=V(|\Phi|)-F(b)$, and choose the potential $V(|\Phi|)=\frac{\lambda}{2}|\Phi|^{4}$, describing short-range repulsive interactions, while relegating the study of more general scenarios to future work.
@@
Furthermore, we have coupled the theory to the metric and gauge field rendering it covariant under general transformations of spacetime coordinates and local $\text{U}(1)_{\text{Q}}$ transformations. Despite being Poincaré invariant the action (48) contains only a single time derivative, reminiscent of nonrelativistic field theories. As a consequence, $\Phi$ and $\Phi^{\dagger}$ are not independent, so they account for only a single propagating degree of freedom [^9] In other words, there is no Higgs mode.. With this in mind, it is convenient to employ the polar decomposition $\Phi=\sqrt{\rho}e^{i\psi}$ such that the action (48) takes the following form
@@ 3.5 Linearized hydrodynamics
Using (70), the speed of sound can be recast in the conventional form landau1987fluid
@@
where the second equality follows after applying a number of thermodynamic identities (see Herzog:2008he ).
@@
Hence, $\text{SDiff}(\mathbb{R}^{2})$ acts on the transverse component $\pi_{\perp}$, forbidding the kinetic term for this mode (see also Endlich:2010hf ). Similarly, the $\varphi$ Goldstone realizes nonlinearly not only the $\text{U}(1)_{Q}$ symmetry but rather the full chemical shift symmetry
@@ 4 Superfluids
We now move our attention to the superfluid phase and introduce an action principle governing the hydrodynamics of the two-fluid model developed by Tisza and Landau in the context of superfluid helium-4 Tisza:1938ydz ; Landau:1941lul . While complex scalar field theories for zero-temperature superfluids are well established, their finite-temperature counterparts are lacking. An EFT for finite-temperature superfluids, formulated in terms of a set of Goldstone fields, was developed by Alberto Nicolis in Ref. Nicolis:2011cs . The model introduced here can be regarded as a UV completion of Nicolis:2011cs .
@@ 4.2 Constitutive relations
Matching to the superfluid constitutive relations Arean:2023nnn we establish the following identifications
@@
in full agreement with Arean:2023nnn .
@@ 4.3 Linear response and second sound
Eq. (91) bears close resemblance to the quadratic theory Eq. (17) of Ref. Nicolis:2011cs . However, Eq. (91) contains fewer free parameters and provides a more microscopic interpretation to some of them. For example, the coefficient in front of $(\partial_{t}\varphi)^{2}$ is fixed by the strength of interaction $\lambda$, that of $\partial_{t}\pi^{i}\partial_{i}\varphi$ is set by equilibrium charge density $\rho_{0}$, and that of $(\partial_{i}\varphi)^{2}$ by the ratio $\frac{\rho_{0}}{m}$.
Furthermore, Eq. (91) differs structurally from Nicolis:2011cs in that it contains a $\partial_{t}\pi^{i}\partial_{i}\varphi$ coupling, instead of the $\partial_{i}\pi^{i}\partial_{t}\varphi$ term. While these terms are equivalent for smooth configurations up to a total derivative, they differ in the presence of singularities associated with superfluid vortices, i.e. when $\epsilon^{\mu\nu\rho}\partial_{\nu}\partial_{\rho}\varphi\neq 0$. To elucidate the role of superfluid vortices let us analyse the equation of motion for $\pi^{i}$,
@@
One of the modes depends only on the entropy (thermal) sector, while the second one, controlled by the charge density, reproduces the familiar dispersion relation of the Bogoliubov phonon pitaevskii2016bose . Therefore, in the small $\lambda$ limit the two solutions decouple into a charge-dominated Bogoliubov phonon and a thermal (entropy-dominated) mode.

### v2

@@ 1 Introduction
Fluid dynamics is one of the oldest and most established fields of physics, with applicability ranging from microscopic quantum fluids, such as the quark-gluon plasma, all the way to the universe itself. Despite its long and illustrious history, a complete understanding of fluids is still lacking. Some of the most pressing open problems include nonlinear phenomena such as turbulence and the million-dollar problem on the Navier–Stokes equations, but also the behavior of strongly correlated quantum fluids, and questions concerning their existence (or lack thereof) at zero temperature Endlich et al. (2011); Dersy et al. (2024); Cuomo et al. (2024).
@@
From a theoretical standpoint, a more compelling picture is offered by a field-theoretic formulation of hydrodynamics in terms of a local action Son (2002); Dubovsky et al. (2006, 2012); Nicolis (2011); Kovtun et al. (2014); Crossley et al. (2017). In this framework, the hydrodynamic equations follow from the Euler–Lagrange equations for the relevant low-energy fields, while the path-integral formalism grants access to the correlation functions and scattering amplitudes (see, e.g., Son (2005); Endlich et al. (2011); Nicolis (2011); Dersy et al. (2024); Jain and Kovtun (2024); Delacretaz and Mishra (2024); Cuomo et al. (2024)). In short, the action formulation allows the use of the full quantum field theory (QFT) toolkit. The models in question are effective field theories (EFTs) where infrared degrees of freedom are the Nambu-Goldstone bosons of spontaneously broken symmetries. Unfortunately, without microscopic input from the UV, these EFTs can suffer from various pathologies.
@@
where $D_{\mu}\psi=\partial_{\mu}\psi-A_{\mu}$ is a covariant derivative and $\mathcal{L}(X)$ is an arbitrary function. For example, choosing $\mathcal{L}(X)=X$ gives a simple quadratic theory with linear equations of motion. In general, however, the resulting equations are nonlinear and lead to the formation of caustic singularities in finite time, wherein the characteristics intersect and second-order derivatives diverge, signaling the breakdown of the EFT Babichev (2016). These issues can be circumvented by introducing a UV completion in terms of a complex scalar field, which reproduces the same low-energy dynamics as the EFT model (up to the formation of caustic singularities in the EFT) while regularizing the caustic instability Babichev and Ramazanov (2017).
For superfluids, a suitable UV completion is given by a canonical complex scalar field action Babichev et al. (2018)
@@
where $D_{\mu}\Phi=\left(\partial_{\mu}-iA_{\mu}\right)\Phi$. By placing the system at finite chemical potential via a constant background gauge field $A_{\mu}=(\mu_{0},0)$ with $\mu_{0}>m$, and implementing the polar decomposition $\Phi=\sqrt{\rho}e^{\psi}$, the theory Eq. (2) reproduces Eq. (1) in the long-wavelength limit after integrating out the Higgs mode Joyce et al. (2022). Clearly, the theory Eq. (2) contains more information than the EFT Eq. (1). For example, it can be used to compute the healing length of the superfluid. It is precisely the absence of information about the short-range physics in the EFT Eq. (1) that manifests itself in the appearance of caustic singularities.
While the UV completion of the superfluid EFT at zero temperature is well understood, the situation becomes more subtle at finite temperatures, where thermal excitations populate the normal component and give rise to a two-fluid description consisting of a (quantum) superfluid coupled to a normal (classical) fluid. An analogous question can be posed regarding the UV completion of charged fluids in the normal phase, wherein the charge and thermal sectors are likewise intertwined.
Motivated by these considerations, in this manuscript, we propose a UV completion to the EFTs of charged fluids Dubovsky et al. (2012) and superfluids at finite-temperature Nicolis (2011). For this purpose, we utilize a geometric picture of fluid flow in terms of a comoving spatial hypersurface and define on it a complex scalar field transforming in the linear representation of the $U(1)_{Q}$ charge symmetry. The models proposed herein reproduce the established constitutive relations for the ideal part of hydrodynamic currents, and the hydrodynamic excitations are identified with the Nambu-Goldstone modes corresponding to small fluctuations around the state of thermodynamic equilibrium.
We show that ordinary charged fluids are necessarily invariant under a restrictive chemical shift symmetry that freezes the dynamics of the elementary charges in the comoving space. In this interpretation, charges exhibit fractonic immobility on the comoving plane but are dragged through spacetime by the fluid flow. This symmetry is relaxed in the superfluid phase in order to account for the dynamics of the superfluid component. Our formulation naturally accommodates a classification of exotic fluid phases interpolating between normal fluids and superfluids. These fracton fluids exhibit partially restricted mobility, a consequence of the comoving multipole symmetries [^1] Analogous ”crystal-multipole” symmetries were introduced in Jain (2024)., which in our approach emerge as a finite truncation of the chemical shift symmetries.
@@ 2 Comoving formulation of fluid dynamics
In this section we review the comoving formulation of perfect fluids and encode their relativistic hydrodynamics in an action principle. The variational approach to fluid mechanics in terms of a set of comoving coordinates has a long history Eckart (1960); Salmon (1988); Bahcall and Susskind (1991), see also Jackiw (2002) for a pedagogical introduction. More recently, the problem has been revisited from a modern effective field theory viewpoint Dubovsky et al. (2006), which we will mostly follow in this section. We also provide a complementary geometrical interpretation of the comoving formulation in terms of spacelike hypersurface embedded in the physical spacetime. Our setup aligns closely with the geometric framework of elasticity theory introduced in Armas and Jain (2020). Although the comoving hypersurface perspective may currently appear as a mere reinterpretation, it sets the stage for the constructions developed in the upcoming sections.
@@ 2.3 Effective action for uncharged fluids
Therefore, the leading order theory for the hydrodynamics of uncharged fluids can be expressed as Dubovsky et al. (2006)
@@
There are also conservation laws associated with $\text{SDiff}(\mathbb{R}^{2})$ symmetry Eq. (6), corresponding to vorticity conservation (see Cuomo et al. (2024)).
Finally, it is important to emphasize that the comoving formulation presented herein does not describe the most general hydrodynamic flows, but rather a subclass of homoentropic flows maintaining constant entropy per fluid element at all times. This is already apparent from the fact that the entropy density, Eq. (20), is fully determined by the mechanical displacements of the fluid elements, so that there is no room for genuine independent thermal transport. In particular, we have formulated the comoving theory in terms of only $d=2$ scalar fields $\phi^{I}$, whereas, in general, hydrodynamics admits $2+1$ independent variables [^5] At the level of linearized ideal hydrodynamics, however, temperature fluctuations do not constitute an independent hydrodynamic mode and can be solved for in terms of velocity gradients via energy conservation.. One way to remedy this is by introducing an “internal clock” variable, thereby enlarging the comoving manifold to 2+12+1–dimensions parameterized by the coordinates $\phi^{M}$ with $M=0,1,2$ Nickel and Son (2011). However, such a choice leads to a generalized Euler relation whose physical interpretation is not clear at the moment Mancilla (2025). Therefore, for the remainder of this work, we will stick with the comoving formulation $\phi^{I}$ with $I=1,2$.
@@ 3 Charged fluids
Utilizing the comoving hypersurface interpretation, we propose a field-theoretic model for charged fluid dynamics that reproduces the nonlinear hydrodynamic equations of a perfect fluid. Contrary to previous formulations Dubovsky et al. (2012), the charge symmetry is realized linearly on the complex scalar field $\Phi$ residing on the comoving hypersurface. We provide a first-principles proof for the emergence of the chemical shift symmetry in charged fluids, corresponding to the comoving conservation of all multipole moments, and elucidate the resulting symmetry structure. Then, we formulate a symmetry-invariant action, derive the constitutive relations, and establish an explicit dictionary mapping the fields of the theory to hydrodynamic variables. We also discuss the equilibrium state, its symmetry-breaking pattern, and the spectrum of Nambu–Goldstone modes.
@@ 3.1 Multipole symmetries on the comoving plane
These charges reflect the conservation of the multipole moments of the charge density, familiar in the context of fractons Gromov (2019), with respect to the comoving hypersurface. Such conservation laws impose strong constraints on the dynamics of the elementary charges giving rise to fractonic behavior, wherein charges cannot move along the comoving hypersurface without violating one of the higher-moment conservation laws [^6] Nevertheless, they can still be advected through spacetime by a fluid’s motion.. A theory of charged fluids must respect the conservation of the multipole moments Eq. (38) and therefore be invariant under the symmetries generated by these charges,
@@
where $\Lambda_{I_{1}\dots I_{N}}$ is a constant symmetric tensor of parameters. Since this must hold for any $N$, we conclude that in ordinary fluids the charge symmetry Eq. (34) must be promoted to the chemical shift symmetry [^7] The chemical shift symmetry was first introduced in Dubovsky et al. (2012) to construct an effective field theory for hydrodynamics. In our formulation, however, it is realized linearly on the complex scalar field and emerges naturally as a consequence of Noether’s theorem.
@@ 3.2 The model
Having established the appropriate degrees of freedom for charged fluids and their symmetry structure, we are now equipped to write down an action governing their hydrodynamics. For this purpose, we consider the following action [^8] The theory Eq. (48) can be understood as a UV completion of the effective hydrodynamic theory constructed in Dubovsky et al. (2012).
@@
where $D_{0}\Phi=u^{\mu}(\partial_{\mu}-iA_{\mu})\Phi$ and $E(|\Phi|,b)$ is an arbitrary function. Throughout this manuscript, we assume that $E(|\Phi|,b)$ is factorizable [^9] This choice corresponds to a subclass of fluids that admit a factorizable equation of state., $E(|\Phi|,b)=V(|\Phi|)-F(b)$, and choose the potential $V(|\Phi|)=\frac{\lambda}{2}|\Phi|^{4}$, describing short-range repulsive interactions, while relegating the study of more general scenarios to future work [^10] In full generality, the action can also exhibit arbitrary dependence on the symmetry-invariant combination $Y\equiv i\left(\Phi^{\dagger}D_{0}\Phi-\Phi D_{0}\Phi^{\dagger}\right)$..
@@
Furthermore, we have coupled the theory to the metric and gauge field rendering it covariant under general transformations of spacetime coordinates and local $\text{U}(1)_{\text{Q}}$ transformations. Despite being Poincaré invariant the action (48) contains only a single time derivative, reminiscent of nonrelativistic field theories. As a consequence, $\Phi$ and $\Phi^{\dagger}$ are not independent, so they account for only a single propagating degree of freedom. With this in mind, it is convenient to employ the polar decomposition $\Phi=\sqrt{\rho}e^{i\psi}$ such that the action (48) takes the following form
@@ 3.5 Linearized hydrodynamics
Using (70), the speed of sound can be recast in the conventional form Landau and Lifshitz (1987)
@@
where the second equality follows after applying a number of thermodynamic identities (see Herzog et al. (2009)).
@@
Hence, $\text{SDiff}(\mathbb{R}^{2})$ acts on the transverse component $\pi_{\perp}$, forbidding the kinetic term for this mode (see also Endlich et al. (2011)). Similarly, the $\varphi$ Goldstone realizes nonlinearly not only the $\text{U}(1)_{Q}$ symmetry but rather the full chemical shift symmetry
@@ 4 Superfluids
We now move our attention to the superfluid phase and introduce an action principle governing the hydrodynamics of the two-fluid model developed by Tisza and Landau in the context of superfluid helium-4 Tisza (1938); Landau (1941). While complex scalar field theories for zero-temperature superfluids are well established, their finite-temperature counterparts are lacking. An EFT for finite-temperature superfluids, formulated in terms of a set of Goldstone fields, was developed by Alberto Nicolis in Ref. Nicolis (2011). The model introduced here can be regarded as a UV completion of Nicolis (2011).
@@ 4.2 Constitutive relations
Matching to the superfluid constitutive relations Areán et al. (2024) we establish the following identifications
@@
in full agreement with Areán et al. (2024).
@@ 4.3 Linear response and second sound
Eq. (91) bears close resemblance to the quadratic theory Eq. (17) of Ref. Nicolis (2011). However, Eq. (91) contains fewer free parameters and provides a more microscopic interpretation to some of them. For example, the coefficient in front of $(\partial_{t}\varphi)^{2}$ is fixed by the strength of interaction $\lambda$, that of $\partial_{t}\pi^{i}\partial_{i}\varphi$ is set by equilibrium charge density $\rho_{0}$, and that of $(\partial_{i}\varphi)^{2}$ by the ratio $\frac{\rho_{0}}{m}$.
Furthermore, Eq. (91) differs structurally from Nicolis (2011) in that it contains a $\partial_{t}\pi^{i}\partial_{i}\varphi$ coupling, instead of the $\partial_{i}\pi^{i}\partial_{t}\varphi$ term. While these terms are equivalent for smooth configurations up to a total derivative, they differ in the presence of singularities associated with superfluid vortices, i.e. when $\epsilon^{\mu\nu\rho}\partial_{\nu}\partial_{\rho}\varphi\neq 0$. To elucidate the role of superfluid vortices let us analyse the equation of motion for $\pi^{i}$,
@@
One of the modes depends only on the entropy (thermal) sector, while the second one, controlled by the charge density, reproduces the familiar dispersion relation of the Bogoliubov phonon Pitaevskii and Stringari (2016). Therefore, in the small $\lambda$ limit the two solutions decouple into a charge-dominated Bogoliubov phonon and a thermal (entropy-dominated) mode.